### v1

# Adacc: Adaptive Compression and Activation Checkpointing for LLM Memory Management

Abstract.
Training large language models often employs recomputation to alleviate memory pressure, which can introduce up to 30% overhead in real-world scenarios. In this paper, we propose Adacc, a novel memory management framework that combines adaptive compression and activation checkpointing to reduce GPU memory footprint. It comprises three modules: (1) We design layer-specific compression algorithms that account for outliers in LLM tensors, instead of directly quantizing floats from FP16 to INT4, to ensure model accuracy. (2) We propose an optimal scheduling policy that employs MILP to determine the best memory optimization for each tensor. (3) To accommodate changes in training tensors, we introduce an adaptive policy evolution mechanism that adjusts the policy during training to enhance throughput. Experimental results show that Adacc can accelerate the LLM training by 1.01 to 1.37 compared to state-of-the-art frameworks, while maintaining comparable model accuracy to the Baseline.
1. Introduction
Large Language Models (LLM) have achieved unprecedented success across various domains. Scaling laws show that model size is crucial for performance, leading to increasingly larger models. For instance, the size of large models has increased by over 360×, from GPT-2 (Radford et al., 2019) with 1.5 billion parameters to PaLM with 540 billion parameters (Chowdhery et al., 2023). This trend is expected to continue, with model sizes exceeding the memory capacity of a single GPU (tens of gigabytes).
To address GPU memory limitations for training large models, the recomputation approach (activation checkpointing) has been proposed. It alleviates memory pressure by discarding activations during forward propagation and regenerating them on demand during backward propagation (Chen et al., 2016). This mainstream method is widely used in frameworks like Megatron (NVIDIA, 2024b) and MindSpore (Huawei, 2024), which implement various policies to manage which tensors are retained or recomputed.
Limitations of existing works. The existing recomputation methods can be placed into two categories. First, coarse-granularity recomputation selects specific tensors for recomputation based on predefined rules, following an all-or-nothing pattern. For instance, Megatron libraries offer a full recomputation option that caches only the inputs to transformer layers as checkpoints, discarding other activations and recomputing them during backward propagation. However, this approach introduces significant recomputation overhead. Specifically, we profile Megatron-LM using a 345M parameter GPT model with an 8 batch size on 4 Tesla V100 GPUs. Our results indicate that full recomputation accounts for over 28% of the total training time. This inefficiency arises because the rule-based policy does not account for the varying memory demands and computational costs of different model layers, as described in §2.2
Second, model-adaptive recomputation methods (Chen et al., 2016, 2024) have been proposed to employ search algorithms to identify the optimal policy based on the characteristics of deep learning models and the capabilities of the devices. For example, Checkmate uses MILP to find a suitable recomputation policy through offline searching. However, due to the vast search space, it is unable to find the optimal solution for the large language model.
| Tensor | Size (MB) | Recmp. time (ms) | Compression time (ms) | Compression overhead (MB) | Recmp. Comp. |
| T1 | 96 | 0.36 | 0.37 | 24 | |
| T2 | 42 | 1.02 | 0.16 | 11.8 | |
| T3 | 42 | 0.58 | 0.16 | 11.8 | |
| T4 | 10.5 | 0.04 | 0.04 | 2.6 |
*Recmp. time: the time recomputed from the previous layer to this layer. *As recomputation frees tensors from the GPU, the space overhead is zero.
New observations. Data compression techniques have been proposed to reduce data size. Surprisingly, advancements in GPU parallel computing and algorithm optimization have significantly improved computation efficiency, making data compression feasible for AI training. We explore compression opportunities by selecting activation tensors from the GPT model and presenting the computational time and memory costs for recomputation and the proposed compression methods (in §3.2). As shown in Table 1, We identify four tensors, each suited to different memory optimization techniques. Specifically, we define the optimization bandwidths, and , to demonstrate the efficiency of each technique. The bandwidth is calculated as . For example, the and for T1 is 266 and 194, respectively. It is better to optimize T1 by recomputation. In contrast, compression is more efficient for T2. Therefore, different memory optimization strategies should be applied to tensors to maximize model training throughput rather than only considering the recomputation as implemented in mainstream frameworks (NVIDIA, 2024b).
Our work. This finding motivates the proposal of a new adaptive memory management framework, Adacc, that integrates compression and recomputation to maximize training throughput and prevent GPU out-of-memory issues. However, Adacc faces three challenges when introducing compression into the training framework. First, directly applying compression into framework can lead to greater loss of model accuracy. To address this, Adacc designs four layer-specific compression algorithms that account for data outliers, offering fast, memory-efficient solutions with minimal impact on training accuracy. Second, how to find the optimal combined policy for compression and recomputation within an acceptable time? To solve it, Adacc uses a custom mixed-integer linear programming (MILP) algorithm to achieve a globally optimized tensor policy for LLM models, significantly reducing the search space by leveraging the repeated training structures in LLMs. Third, the dynamic variation of data outliers during training affects the model’s compression rate, rendering the current policy suboptimal. Therefore, Adacc adaptively adjusts the model policy efficiently in response to data changes to enhance training efficiency.
We evaluate Adacc on a wide range of benchmarks and downstream tasks across diverse scenarios, achieving significant performance gains without compromising model accuracy.
2. Background and Motivation
2.1. Large language model training
Large language models, composed of multiple repeated layers (e.g., Transformer block), are trained iteratively with data batches. Each step involves forward propagation (FP) to compute large activations and backward propagation (BP) to calculate gradients for optimization. A model with M parameters requires 16M + M bytes of memory, including 2M FP16 parameters, 2M FP16 gradients, 12M FP32 optimizer data (e.g., Adam), and M FP16 intermediate activation, where depends on the user-defined batch size. Generally, training samples are typically processed in large batches to enhance throughput and device utilization by increasing arithmetic density (Chen et al., 2024), leading to substantial GPU memory usage for activations.
However, GPUs’ limited memory capacity restricts large model training, as both model states and activations (feature maps) consume memory. Model states consist of parameters, gradients, and optimizer states, such as momentum and variance in algorithms like Adam. Activations are intermediate data generated during training that need to be stored on the GPU, with memory consumption increasing as batch size grows. Users often use large training batches to maximize GPU utilization, which exacerbates memory bottleneck issue (i.e., memory wall). For example, in GPT model training, 66% of memory is consumed by stashed immediately activations on 4 V100 GPUs with a batch size of 4. This number grows to 76% for 8 batch size. Therefore, optimizing activations is essential for efficient LLM training.
2.2. Existing works and their shortcomings
Several works have explored optimizations for GPU memory wall issue. Prior wroks use activation checkpointing (recomputation) to extend the limited capacity of GPU memory (Chen et al., 2016; Jain et al., 2019; Chen et al., 2024). However, we find not all activations are suited for recomputation. Small activations that are costly to recompute should remain in GPU memory or be optimized with other techniques, as tensors within the red circle shown in Figure 1. Some works propose lossless and lossy encodings to reduce activation memory (Jain et al., 2018; Rhu et al., 2018), but these target traditional CV models, whereas Adacc focuses on large language models. Some methods swap activations from GPU to CPU to save memory (Rhu et al., 2016; Chen et al., 2022), but they suffer performance loss due to limited PCIe bandwidth and require intrusive framework modifications, making them impractical. Additionally, intra- and inter-operator parallelism methods split models and distribute partitions across GPUs to reduce memory footprints (Rajbhandari et al., 2020; Shoeybi et al., 2019; Huang et al., 2019), complementing Adacc.
3. Design
Adacc aims to train LLMs on one or more GPUs, when GPU memory or a user-specified budget is insufficient. It keeps the training parameters, such as batch size, unchanged while optimizing performance and preserving accuracy. At the high level, Adacc adopts the idea of adaptive activation checkpointing and compression. While, it introduces three challenges. (1) How to design compression algorithms that minimize accuracy loss? (2) How to optimize memory usage with compression and recomputation? (3) How to dynamically adjust policies during training to maximize throughput?
3.1. Overview
Adacc consists of three key modules: the model profiler, policy maker, and model rectifier, which collaborate to achieve the design goals. Figure 2 provides an overview of Adacc.
Model profiler. Before deploying a model, a test run is conducted with user-defined configurations (GPU features, model architecture, batch size, etc.). Metrics like execution time, operator size, and dependencies are recorded in a database for the policy maker’s scheduling decisions ❶. Based on the profiled tensor size and our defined compression algorithm, we calculate the compression rate and the time cost for both compression and decompression ❷.
Policy maker. Next, we input the collected operator size, computation time, tensor compression rate, and compression/decompression time into the policy maker’s MILP cost model ❸ and determine the optimization policy for each tensor ❹. The training framework then implements the optimal schedule defined by the model policy maker, deploying the model for training on physical devices using the training framework ❺.
Policy rectifier. The monitor periodically collects data characteristics during training to detect if the policy is suboptimal and adjusts it promptly ❻.
3.2. Layer-specific Compression
Traditional compressions. A straightforward approach is to compress activations from FP16 to INT4 using per-token quantizaiton (as shown in Figure 3(b)(i)). However, our experience shows that this naive approach can reduce model accuracy by up to 39% in GPT models or even leads to gradient overflow in training (as shown in §4.3). To solve this challenge, we design several specific compression algorithm for different LLM layers.
Outlier-separated activation compression. As shown in Figure 3(a), there are many outliers in LLM tensors. Existing methods quantize outliers with other values, resulting in significant outliers accuracy loss and greater model degradation. To address this, a naive method groups activations by channel and quantize outliers separately from the rest of the values, as shown in Figure 3(b)(ii). However, since the data primitively is stored along token dimension (row), data adjacent on the channel dimension (column) is not contiguous in memory. Since the compressor requires contiguous data in memory, we need to transpose and align the tensor before compression. However, performing these operations can be time-consuming. Therefore, we design an outlier-separated method to isolate the outlier from normal value, and quantize the normal layer based on their original alignment, as shown in Figure 3(b)(iii). We use Z-Score method to extract outlier channels from activations, and the Z-Score for each channel is calculated using formulation:
where is the absolute summation of th channel, and represent the average absolute summations, and the standard deviation of absolute summations respectively. We set a Z-Score threshold to extract outlier channels, and it is 3 in our work (experiments show that this setting achieves the best model accuracy). Outlier channels are set to zero in original activation. The extracted outlier channels and their indices are stored without compression to maintain model accuracy, while the normal activations are compressed from FP16 to INT4 to reduce memory usage. The compression and decompression workflows are shown as Figure 3(c).
Activation quantization. We regard tensors of arbitrary dimensions as one-dimensional array stored continuously in memory, and group the activations. For symmetric quantization, in each quantization group, scaling factor is calculated by , where is the max absolute value, and is the max absolute value 4bit signed integer can represented. After that, each value is calculated by formulation:
where means nearest rounding, is the quantized value, and clip ensures the greater than 7 are capped at 7. During dequantization, the activations are restored using formulation: . For majority of activations, which are distributed on both side of zero, symmetric quantizaiton can be applied to accelerate quantizaiton (e.g., T=[-2,-1,0,1,2]).
For asymmetric quantization, we calculate offset for each group by , and the scaling factor is calculated using , where and is the maximum and the minmum value in the group, respectively. Each value is calculated by formulation:
during dequantization, the activation are restored using formulation: . Asymmetric quantizaiton are applied to the activations, which distributes on only one side of zero to minimize quantizaiton errors (e.g., T=[1,2,3]).
Layer-specific compression. We apply different compression scheme to different activations based on their features.
For Linear, LayerNorm, and GELU layers with many outliers, we apply an outlier-separated activation compression scheme to minimize errors.
For the Query, Key, and Value matrixs, we using channel-wise symmetric quantization, as no extra memory alignment is needed.
For the Softmax layer and Score, with activations distributed between (0, 1) and no outliers, we apply asymmetric quantization to save ensure accuracy.
For the Dropout mask in attention layers, we compress the original byte format to bit format without accuracy loss.
3.3. Determining Cost-Effectiveness of Tensor Optimization
| Constant variables: | |
| Number of operators | |
| GPU memory capacity | |
| Fixed memory used for storing static data (e.g., model states, gradients, and optimizers) | |
| Number of transformer layers | |
| Optimization variables: | |
| Boolean variable. If is true, it means will follow this policy | |
| Intermediate variables: | |
| Memory size for | |
| Computation time for | |
| Compression time for | |
| Decompression time for | |
| Compression Rate for | |
As analysis in §2.2, selecting an appropriate memory optimization policy for different tensors is crucial for achieving optimal training efficiency. Each layer can opt for either layer-specific compression or recomputation. To find the policy, existing approaches build cost models for all operators, resulting in a large search space, which is impractical for large language models (Jain et al., 2019). Surprisingly, LLMs are composed of repeated identical structures, such as transformer blocks in GPT, with consistent memory and computation requirements, as well as a similar outlier rate. As a result, the locally optimal recomputation policy for a single layer can be applied to other identical layers without searching the global space.
MILP cost model. We use a mixed-integer linear program (MILP) to formulate and describe the problem. Here, we illustrate the design with a transformer-based model. However, our approach also applies to other large deep learning models.
Problem definition. The DNN model comprises operators () that perform training operation saccording to the model topology. represents the optimization policy for . , , and represent recomputation, compression, and retention in GPUs, respectively. Table 2 summarizes all variables used. The objective is to minimize training time (i.e., minimizing the recomputation, compression and decompression overhead) while avoiding out-of-memory issues, which can be formulated as:
| subject to | Dependency constraints | |||||
| Memory constraints | ||||||
Dependency constraints. We constrain each optimization operation for every operator to be executed only once in Equation 5. Since all recomputations require an initial layer as a checkpoint, we store the output of the first layer of each Transformer as the checkpoint. Thus, the output of is either retained or compressed in the GPU, as shown in Equation 6.
Memory constraints. The data stored in the GPU should not exceed the GPU’s memory. GPU memory mainly consists primarily of static memory and activations, so it’s crucial to ensure them stays within the GPU’s capacity, as shown in Equation 7. Activation consists of tensors residing GPU memory () and remaining tensors after compression (), as shown in Equation 8.
Solving MILP. We generate the recomputation policy by using PulP (Python, 2023). Our design drastically reduces the search space compared to existing approaches that model all operators, taking less than 0.5 second to find an optimal policy, as a negligible time relative to the overall training time.
3.4. Adaptive Policy Evolution
As shown in Figure 4, the outliers in tensors vary across different training stages, causing the compression ratio to change dynamically during the training process. Therefore, we must track outliers changes to update memory optimization policy and maximize training throughput. While, tracking outliers requires traversing the entire tensor, making it time-consuming. This introduces a challenge: high-frequency tracking incurs more overhead but improves policy efficiency, while low-frequency tracking minimizes overhead but leads to suboptimal training throughput. Therefore, how to design an optimal tracking frequency is important.
Determining the tracking frequency. We have a new observation that the number of outliers fluctuates significantly during the early training period but stabilizes in the later stages, as shown in Figure 4. Thus, we design to set a high tracking frequency in the early training stage and lower it in the later stages. In Adacc, we apply the Exponential Backoff Algorithm to determine the tracking interval (e.g., iteration), minimizing tracking overhead. Thus, we label the training iterations as either normal iterations or tracking iterations.
Policy evolution mechanism. During training, at each tracking iteration, the framework re-evaluates outliers in the Linear, LayerNorm, and GELU layers and inputs the updated tensor information (e.g., ) into the MILP cost model. If the memory optimization policy changes, the framework adjusts it and continues training with the updated policy.
4. Evaluation
4.1. Evaluation Setup
Platform. We evaluate Adacc on a V100 node, which is equipped with 256GB DRAM, 2 Intel Xeon Gold 6130 CPUs, and 8 32GB Tesla V100 GPUs interconnected via NVLink.
Baseline. We compare Lynx with Megatron-LM (NVIDIA, 2024a), Quantization, and Baseline (NVIDIA, 2024b). (1) Megatron-LM, a widely used framework for large model training, supports full recomputation, checkpointing only transformer layer inputs and discarding other activations. (2) Quantization compresses data directly from FP16 to INT4. (3) Baseline refers to training without any memory optimization, maintaining full model accuracy. (4) Adacc w/o outlier is Adacc with no outlier extracttion. We apply the ZeRO technique to train all the models (Rajbhandari et al., 2020).
4.2. General Result
| Model | Baseline | Full Recomp. | Quantization | Adacc |
| GPT-117M | 57 | 136 | 98 | 136 |
| GPT-345M | 8 | 61 | 19 | 61 |
| GPT-4.7B | 1 | 4 | 3 | 4 |
Training throughput. To evaluate the effectiveness of Adacc, we measure the DNN training throughput across different memory optimization approaches on two small GPT models and one large GPT model, as shown in Figure 5. There are several observations. (1) Adacc has the best training performance among counterparts. Specifically, Adacc improves the throughput by 1.01-1.37 and 1.09-1.28 compared to Full Recomputation and Quantization. (2) Adacc adaptively selects the optimal strategy based on memory requirements to maximize training throughput. For example, with a small batch size, GPU memory is sufficient, and Adacc introduces no additional overhead. In contrast, Full Recomputation and Quantization only provide open and closed options, potentially leading to excessive memory optimization. (3) For GPT-345M, as the batch size increases (e.g., from 1 to 15), the throughput of Adacc initially rises and then begins to decrease. The performance gains from increased GPU utilization with larger batch training outweigh the overhead introduced by compression and recomputation. (4) As the batch size increases, leading to higher GPU memory demands, Adacc tends to recompute more tensors (the policy approaches Full Recomputation) while consistently maintaining higher throughput than Full Recomputation.
Memory footprint reduction. We use the batch size to represent the degree of memory footprint reduction. Table 3 presents the maximum batch size for different counterparts. We observe that Adacc always achieves the maximum batch size. Compared to Baseline and Quantization, Adacc promotes the maximum batch size by up to 2.38-7.62 and 1.38-3.2. This is because Baseline trains models without optimization techniques, while Quantization incurs significant space overhead by storing compressed data for all layers.
4.3. Accuracy
We explore the impact of Adacc’s design on model accuracy, presenting both loss values and downstream task accuracy.
Evaluations on validation loss. We show the loss curve of validation dataset on GPT-117M and GPT-345M across different batch sizes. Figure 6 shows that Adacc achieves comparable loss to the baseline, with only a 0.46%-0.5% difference in loss across two models. Moreover, Quantization performs well in the early stages of training but fails to converge in the later stages due to significant data accuracy loss.
MODEL BS METHOD Lambada Arc-e Sciq Swag Ave. 117M 32 Baseline 53.46 39.56 71.80 37.46 50.57 Quantization 2.77 25.08 - 24.66 13.13 Adacc 53.13 39.1 70.80 37.22 50.06 Adacc w/o outlier 53.03 39.18 69.90 36.98 49.77 64 Baseline 54.76 42.05 72.60 38.19 51.90 Quantization 3.14 25.08 - 24.66 13.22 Adacc 54.82 41.04 71.80 38.09 51.44 Adacc w/o outlier 54.7 39.77 71.80 38.09 51.03 345M 8 Baseline 56.97 42.85 72.50 39.49 52.95 Quantization 43.1 25.08 - 24.66 23.21 Adacc 56.29 42.55 72.70 39.26 52.7 Adacc w/o outlier 56.06 41.79 71.60 39.12 52.14 16 Baseline 58.06 44.65 76.20 40.88 54.95 Quantization 50.45 25.08 - 24.65 25.05 Adacc 57.67 44.53 76.60 40.79 54.90 Adacc w/o outlier 57.32 43,77 75.10 40.17 54.09
Evaluations on downstream tasks. We evaluate zero-shot downstream tasks of our pretrained GPT-117M and 345M models. Each models is trained using two different batch sizes respectively, and applied different memory optimization methods to observe the accuracy of our pre-trained models on zero-shot tasks. As shown in Table 4, our method has less than 0.5% accuracy degradation compared to the Baseline (oracular method), while Quantization has average 39% accuracy loss.
4.4. Ablation Study
The effectiveness on layer-specific compression. We investigate the efficiency of layer-specific compression in Adacc. As shown in Figure 6, our layer-specific compression achieves better convergence efficiency compared to Adacc without outlier compression. Specifically, layer-specific compression reduces the loss by 0.2% and 0.31% on GPT-117M for batch sizes of 32 and 64, respectively. For the larger GPT-345M, Adacc reduces the loss by 0.48% and 0.7% for batch sizes of 8 and 16, respectively.
The effectiveness on adaptive policy evolution. We evaluate Adacc with and without adaptive policy evolution mechanism. We use two models with 117M and 345M parameters across four batch sizes. Figure 7 illustrates that the throughput with the Adacc adaptive policy is increased by 1.13-1.18 and 1.11-1.2 on GPT-117M and GPT-345M, respectively.
5. Conclusion
In this paper, we propose the Adacc framework for large language model training with data compression and activation checkpointing memory optimization techniques. First, we design layer-specific compression algorithms that account for outliers in LLM tensors, ensuring high training accuracy. Second, we propose an MILP-based scheduling policy to optimize memory for each tensor. Third, we introduce an adaptive policy evolution mechanism that adjusts policies during training to improve throughput. The results show that Adacc outperforms the SOTA method by up to 1.37.